# Federated Unlearning Made Practical: Seamless Integration via Negated Pseudo-Gradients

### v1

@@ V-A Datasets, Models, and Learning Setting
Model Dataset Clients Data Heterogeneity Pre-trained $E$ Task # Unlearning Clients ResNet-18 CIFAR-10 10 LDA ($\alpha$=0.3) $\times$ 1 $\mathcal{C}$ Up to one ResNet-18 CIFAR-100 10 LDA ($\alpha$=0.1) or IID $\times$ 1 or 10 $\mathcal{C}$ Up to two MiT-B0 CIFAR-100 10 LDA ($\alpha$=0.1) 1 $\mathcal{C}$ Up to two ResUNET ProstateMRI 10 Real-world feature skew $\times$ 1 $\mathcal{S}$ Up to one
TABLE II: Settings of reported results. $\mathcal{C}$=classification task, $\mathcal{S}$=segmentation task. $E$=local epochs during FedAvg. # Unlearning Clients reports if the experiments considers a single client or multiple clients to unlearn.
@@ V-B Baselines
We compared PUF against two state-of-the-art baseline methods, which are PGA [18], and NoT [19]. These methods were selected because, similarly to PUF, they do not depend on impractical assumptions, such as access to supplementary data or historical client updates (as we detail in Section II). In fact, requiring a complete history of client updates would necessitate linking each client to their previous submissions, thereby compromising the privacy-centric design of FL. Similarly, assuming the availability of useful public data on the server side to restore the performance of the unlearned model is a strong assumption that is often unrealistic in practice.
Since PGA requires full client participation, we ensured that the original model was trained with full client participation to guarantee a fair comparison. In some experiments, we included a Natural baseline, which bypasses any unlearning procedure before the recovery phase. This allows us to assess whether the influence of the unlearning client would naturally vanish over time without intervention.
@@ V-C Metrics
We evaluated the methods based on two key aspects: their efficiency in recovering performance and their effectiveness in achieving unlearning. When evaluating the effectiveness of unlearning, the primary criterion is minimizing the discrepancy with the metrics of the gold-standard retrained model. This means that neither higher nor lower metric values are inherently preferable; rather, the objective is to reduce the absolute difference between the metrics of the approximate unlearning method and those of the retrained model. We use $\Delta$ as a prefix to denote the absolute difference from the retrained model’s value for a specific metric (e.g., $\Delta$ Test Accuracy represents the absolute difference in Test Accuracy).
@@
Rounds to Recover Performance (Recovery Rounds). We measured the number of FL rounds required by each evaluated method to recover or exceed the test accuracy of the corresponding retrained model. We refer to such a recovery period as recovery phase. Lower values indicate faster recovery, which is preferable.
Communication Efficiency (CE) over retraining. We also report the communication reduction relative to producing a retrained model. Since none of the baselines we consider introduce additional per-round communication overhead compared to FedAvg, we compute CE as the number of rounds needed to learn a retrained model over the recovery rounds required by the specific FU algorithm.
Setting Algorithm Rounds ($\downarrow$) CE ($\uparrow$) Test Acc. Forget Acc. MIA [38] MIA [39] CIFAR-100, IID, $E$=1, ResNet-18 Original 59.91±0.0 79.82±0.7 78.23±0.6 71.42 ±0.6 Retrain 58.32±0.5 58.01±0.7 55.62±0.7 48.86 ±0.6 PGA [18] 12.1 ±4.2 16.7$\times$ 59.12±0.6 72.81 (14.81±1.0) 71.83 (16.15±1.8) 63.02 (14.12±1.6) NoT [19] 6.51 ±1.1 30.8$\times$ 58.94 ±0.4 70.51 (12.61±0.8) 69.31 (13.72±1.9) 63.34 (14.47±0.8) PUF-Special 4.1±1.0 48.8$\times$ 58.82 ±0.3 62.84 (4.86±1.08) 61.49 (5.88±1.42) 55.82 (6.98±1.13) PUF-Regular 5.2±0.63 38.5$\times$ 58.78±0.4 62.36 (4.39±1.01) 60.83 (5.23±0.92) 55.52 (6.68±1.07) CIFAR-100, Non-IID, $E$=1, ResNet-18 Original 53.81±0.0 62.93±6.9 75.44±7.0 61.04 ±7.8 Retrain 51.02±1.4 33.52±4.5 44.02±5.5 32.05 ±5.2 PGA [18] 7.0±5.1 28.6$\times$ 51.42 ±0.6 37.43 (4.32±4.1) 49.21 (6.44±4.90) 36.62 (5.33±5.3) NoT [19] 12.8±6.8 15.6$\times$ 51.31 ±0.2 43.31 (9.9±4.4) 55.75 (11.7±5.7) 44.32 (12.3±5.5) PUF-Special 7.3±4.14 27.4$\times$ 51.49±0.4 35.04 (2.42±2.04) 46.74 (3.66±2.49) 34.34 (2.72±2.22) PUF-Regular 8.0±4.16 25.0$\times$ 51.28±1.46 33.74 (2.37±1.76) 45.51 (3.07±2.42) 33.09 (2.14±2.08) CIFAR-100, Non-IID, $E$=10, ResNet-18 Original 50.32±0.0 60.81±6.63 73.42±6.6 61.31 ±7.5 Retrain 48.02±0.8 31.61±4.5 42.12±5.2 31.24 ±4.9 PGA [18] 4.4±3.8 45.5$\times$ 49.02±1.0 37.34 (6.27±8.1) 49.65 (7.7±9.6) 37.72 (6.6±8.2) NoT [19] 7.7±1.49 26.0$\times$ 48.71±0.7 42.13 (10.5±3.3) 55.72 (13.6±3.3) 45.01 (13.8±4.2) PUF-Special 5.4±1.26 37.3$\times$ 48.36±0.75 32.75 (3.11±2.22) 44.71 (3.71±2.6) 33.92 (3.99±2.28) PUF-Regular 6.3±1.25 31.8$\times$ 48.53 ±0.88 32.48 (1.88±1.6) 43.97 (2.03±1.88) 33.17 (2.44±1.8) CIFAR-10, Non-IID, $E$=1, ResNet-18 Original 83.72 ±0.02 88.64 ±3.93 84.41 ±5.42 81.41 ±6.33 Retrained 83.52 ±1.64 81.11 ±8.04 73.16 ±13.71 72.52 ±10.82 PGA [18] 11.5±5.6 17.4$\times$ 84.01 ±1.0 84.31 (3.20±4.9) 78.54 (5.40±7.7) 77.02 (4.51±6.0) NoT [19] 16.9±8.6 11.8$\times$ 83.81 ±1.2 84.31 (3.10±3.0) 78.04 (4.91±4.5) 75.33 (2.90±3.4) PUF-Special 10.4±6.0 19.2$\times$ 83.81 ±0.2 82.32 (2.01±1.96) 73.33 (3.21±2.72) 73.15 (2.54±2.16) PUF-Regular 13.1±12.1 15.3$\times$ 83.72±1.6 82.22 (1.90±1.5) 73.01 (3.6±3.4) 73.03 (2.52±2.1) CIFAR-100, Non-IID, $E$=1, MiT-B0 Original 75.03 ±0.00 84.25 ±5.63 77.03 ±8.57 74.03 ±6.43 Retrained 73.30 ±0.78 57.80 ±5.29 48.59 ±4.96 44.59 ±3.88 PGA [18] 6.9 ±4.28 7.2$\times$ 73.60 ±0.36 62.15 (4.35±3.28) 53.57 (4.98±3.67) 50.50 (6.91±3.21) NoT [19] 15.5±10.17 3.2$\times$ 73.44±0.74 67.85 (10.05±4.07) 59.42 (10.07±6.59) 53.99 (9.40±5.39) PUF-Special 13.8±5.09 3.6$\times$ 73.55±0.7 60.01 (2.54±1.37) 51.04 (2.46±1.24) 46.71 (2.11±1.11) PUF-Regular 17.0±8.27 2.9$\times$ 73.44±0.76 59.52 (2.00±1.17) 50.61 (1.97±1.18) 47.29 (2.11±1.15)
TABLE III: Performance of PUF and other baselines across different settings, including data distribution and the number of local epochs, with ResNet-18 or MiT-B0. Communication Efficiency (CE column). Baseline results are expressed as mean metric value (mean $\Delta$$\pm$ standard deviation), with $\Delta$ in parentheses representing the average absolute difference from the retrained model, highlighted in blue for better readability. Lower $\Delta$ corresponds to better unlearning performances. Note that, the baselines’ Test Acc. is, by definition, higher than that of the retrained model, as we consider the recovery phase complete once Test Accuracy surpasses that of the retrained model.
@@
(c) CIFAR-100, non-IID, MiT-B0.
Figure 4: Performance comparison of PUF with baselines for different experimental settings indicated in subcaption as a triple dataset, data distribution, model architecture. A smaller polygon represents better unlearning performances. For our method, the charts only visualize the performance of PUF-Special for better visualization. As reported in Table III with the complete results, PUF-Regular performs very similarly to PUF-Special.
(a) CIFAR-100, IID, ResNet-18.
(b) CIFAR-100, Non-IID, ResNet-18.
(c) CIFAR-100, Non-IID, MiT-B0.
Figure 5: Evolution of generalization ability (test accuracy) and forgetting effectiveness (forget accuracy) during the recovery phase across three different settings. Each pair of images presents Test Accuracy (Left) and Forget Accuracy (Right) for a representative client in a specific setting indicated in subcaption as a triple dataset, data distribution, model architecture. For our method, the charts only report the performance of PUF-Special for better visualization.
V-D Main Results
Table III presents the performance of PUF best configurations (hyper-parameter tuning discussed in Sec. V-E) and other baseline methods on different datasets (federated CIFAR-100 and CIFAR-10 datasets), data distributions, and model architectures (ResNet-18, MiT-B0). When not differently stated, the unlearning task considers a single client, and the results are averaged over ten independent experiments, each involving a different client designated as the unlearning client. For results on multiple unlearning (two unlearning clients), the results are averaged over five independent experiments, selecting the set of unlearning clients randomly.
Overall Performance. Figure 4 provides a visual comparison of the performance of PUF-Special against other baselines, where a polygon with a smaller area indicates better overall performance. Forgetting effectiveness is assessed using three metrics: the average absolute gap ($\Delta$) in Forget Accuracy and two variations of MIAs. The efficiency of each FU method is represented by the number of Recovery Rounds required. Additionally, the chart evaluates the FU method’s ability to preserve the global model’s generalization. This is measured by the degradation in Test Accuracy introduced by the unlearning process, quantified as the difference in Test Accuracy between the retrained model and the unlearned model immediately after unlearning is applied — before recovery begins (ideally, at round 0 of recovery).
Overall, PUF demonstrates (1) superior forgetting effectiveness compared to other baselines, and (2) comparable or superior efficiency, except in the CIFAR-100, non-IID, MiT-B0 setting (see Figure 4.c and the 4th row of Table III), where PUF requires a longer recovery period than PGA but achieves significantly better forgetting effectiveness.
Performance Consistency During Recovery. Figure 5 illustrates the evolution of Test Accuracy and Forget Accuracy throughout the recovery phase, beginning at recovery round 0, which corresponds to the application of the unlearning routine (either PGA, NoT, or our proposed solution, PUF). We do not report the evolution of the considered MIAs, as they closely follow the trends of Forget Accuracy.
The figure presents results for three different settings, as indicated in the subcaptions. For reference, we also include values for both the Original Model and the Retrained Model. We emphasize the following key aspects for the full and easy understanding of the reported charts:
@@ V-D Main Results
Thus, Figure 5 provides valuable insights into the dynamics of these metrics during the recovery phase. The analysis highlights that PUF produces a superior unlearned model, significantly narrowing the gap with the retrained model earlier in the recovery phase compared to the other baselines. Additionally, the figure clearly shows that the other baselines fail to achieve accurate forgetting by the end of the recovery phase, as evidenced by their larger gaps in Forget Accuracy.
Per-round Computational and Communication Cost. PUF’s unlearning round has the same computational and communication requirements as a regular FedAvg training round, as both PUF-Special and PUF-Regular achieve effectiveness using the same number of local epochs as standard training.
On the other hand, PGA requires substantially more local epochs to generate the unlearned model—up to 5$\times$—to achieve satisfactory unlearning performance, as reported in its original studies. NoT does not require any client-side computation to produce the unlearning model but exhibits less efficiency in terms of longer recovery phase.
Cumulative Computational and Communication Cost. NoT requires significantly more recovery rounds to restore the performance of the unlearned model, resulting in lower cumulative computational and communication efficiency compared to PUF (see the CE column in tabular results). In the CIFAR-100, Non-IID, MiT-B0 setting, PGA demonstrates greater efficiency than PUF, requiring fewer recovery rounds and incurring lower cumulative costs. However, it produces a less effective global model in terms of forgetting metrics, as already highlighted above in this discussion.
Seamless Integration into FedAvg. PUF-Regular can be effortlessly incorporated into regular FedAvg rounds by reversing the direction of the unlearning clients’ (aggregated) updates. Unlike methods such as PGA, or PUF-Special, PUF-Regular does not alter the regular training process or require a dedicated unlearning round to remove clients’ contributions.

### v2

@@ IV-C Discussion
V Experimental Results
V-A Datasets, Models, and Learning Setting
We conducted a comprehensive set of experiments on image classification tasks and image segmentation tasks. For image classification, we used federated versions of the CIFAR-10 and CIFAR-100 datasets [33], which consist of 60,000 32×32 color images. CIFAR-10 comprises 10 classes, while CIFAR-100 includes 100 classes. The datasets were partitioned to simulate 10 clients, ensuring no overlapping samples among them. The experiments were performed under both Identically and Independently Distributed (IID) and non-IID settings across clients. To create non-IID data, we applied label-skew partitioning based on a distribution determined by Latent Dirichlet Allocation (LDA) [34], using concentration parameters of $\alpha=0.3$ for CIFAR-10 and $\alpha=0.1$ for CIFAR-100. Figure 3 provides a visual depiction of the label distribution among clients. For image classification, we employed a standard ResNet-18 [35] and a visual transformer, namely the MiT-B0 [36]. Before performing unlearning, we train ResNet-18 from scratch for 200 FL rounds while we fine-tune MiT-B0 for 50 FL rounds, starting from a pre-trained checkpoint. If not differently indicated, clients run one local epoch ($E=1$) for standard FedAvg.
We also design a set of experiments for sample unlearning, using the same configurations and hyperparameters as in the client unlearning setup. When evaluating sample unlearning, 50% of the samples in the local dataset of each client that requests to be forgotten are removed. The forget subset is obtained through uniform random selection from the client’s local dataset. The remaining samples (retain data) are then used in the recovery phase for all considered methods, meaning that the target clients contribute to the recovery process only with their retain data.
Table II outlines the settings of our experiments. For each setting, we performed a variable number of experiments, considering one or more clients to forget. In all experiments, we consider SGD as a server-side optimizer with $\eta_{s}$=1.0, and we set $\eta_{r}$=1.0 for PUF-Regular. If not differently indicated, we set $\eta_{u}$=2.0 for PUF-Special and $\eta_{u}$=20.0 for PUF-Regular (Section V-E discusses the tuning of our methods). Further details about per-setting hyper-parameters are reported in the following.
@@ V-B Baselines
We compared PUF against five state-of-the-art baseline methods, which are FedEraser [16], PGA [19], FedAU [21], MoDe [20] and NoT [22]. Our comparisons primarily focus on FU methods that avoid impractical assumptions, such as the need for historical information. Nevertheless, we also include FedEraser [16], despite its requirement to store past client models, as it serves both as a widely adopted reference baseline in the literature and as a means to illustrate the overhead incurred by methods that rely on this requirement. Hyper-parameter tuning of baseline methods can be found in Appendix A (Supplementary Material).
As PGA [19] and NoT [22] were originally evaluated with full client participation, both during the training of the original model and in the recovery phase, we adopted the same setting to ensure a fair comparison.
In some experiments, we included a Natural baseline, which bypasses any unlearning procedure before the recovery phase. This allows us to assess whether the influence of the target client would naturally vanish over time without intervention.
@@ V-C Metrics
We evaluated the proposed methods along two key dimensions: their efficacy in achieving unlearning and their efficiency in recovering model utility. For unlearning efficacy, the primary goal is to minimize the discrepancy between the metrics of the approximate unlearning method and those of the gold-standard retrained model. In this context, neither higher nor lower absolute metric values are inherently preferable; instead, the objective is to reduce the absolute difference with respect to retraining from scratch. We denote this absolute difference with the prefix $\Delta$ (e.g., $\Delta$ Test Accuracy represents the absolute difference in Test Accuracy between the unlearned model and the retrained model). For unlearning efficiency, we measure the cumulative computational (FLOPs), communication (bytes), and storage (bytes) overhead required to perform unlearning and to recover model utility. In addition, we report the relative improvement over retraining from scratch, where higher gains indicate better performance.
@@
Communication, Computational, and Storage Costs. We compute cumulative communication, computational, and storage costs for each federated unlearning method, considering both the unlearning phase and the recovery phase. In the unlearning phase, we quantify these costs according to the original design of each algorithm, under consistent assumptions between methods. The recovery phase corresponds to standard FedAvg rounds involving only retained clients; all methods incur the same per-round costs during recovery, differing only in the number of recovery rounds required to restore model utility. For retraining baselines, the reported costs correspond to training the model from scratch with the remaining clients, i.e., the total FedAvg rounds performed after removing the target clients. We provide a detailed explanation of the overhead calculations for each method in Appendix B (Supplementary Material).
Setting Method Efficacy Efficiency Test Acc. Forget Acc. ($\Delta$ $\downarrow$) MIA[Song] ($\Delta$ $\downarrow$) MIA[Yeom] ($\Delta$ $\downarrow$) Communication Computation Storage Bytes ($\times$ $\uparrow$) FLOPs ($\times$ $\uparrow$) Bytes ($\times$ $\uparrow$) CIFAR-100, IID, ResNet-18, EE=1 Original 59.9±0.0 79.8±0.7 78.2±0.6 71.4±0.6 Retrain 58.3±0.5 58.0±0.7 55.6±0.7 48.8±0.6 1.62e11 (1.0×) 1.35e15 (1.0×) 4.49e07 (1.0×) FedEraser 58.4±0.5 59.5 (1.5±0.7) 57.6 (2.0±0.5) 49.9 (1.1±1.1) 1.62e11 (1.0×) 6.75e14 (2.0×) 8.98e10 (0.0005×) PGA 59.1±0.6 72.8 (14.8±1.0) 71.8 (16.1±1.8) 63.0 (14.1±1.6) 9.86e09 (16.4×) 8.54e13 (15.8×) 4.94e08 (0.0909×) MoDe 58.7±0.5 69.0 (11.1±0.7) 70.2 (14.6±2.0) 62.4 (13.5±0.8) 1.75e10 (9.3×) 1.46e14 (9.3×) 8.98e07 (0.5×) FedAU 59.0±0.5 74.7 (16.8±1.0) 76.3 (20.6±2.1) 61.6 (12.8±2.4) 2.67e09 (60.8×) 2.23e13 (60.6×) 4.49e07 (1.0×) NoT 58.9±0.4 70.5 (12.6±0.8) 69.3 (13.7±1.9) 63.3 (14.4±0.8) 5.25e09 (30.9×) 4.39e13 (30.8×) 4.49e07 (1.0×) PUF-Special 58.8±0.3 62.8 (4.9±1.1) 61.5 (5.9±1.4) 55.8 (7.0±1.1) 3.40e09 (47.6×) 2.84e13 (47.5×) 4.49e07 (1.0×) PUF-Regular 58.8±0.4 62.4 (4.4±1.0) 60.8 (5.2±0.9) 55.5 (6.7±1.1) 5.10e09 (31.8×) 4.26e13 (31.7×) 4.49e07 (1.0×) CIFAR-100, Non-IID, ResNet-18, EE=1 Original 53.8±0.0 62.9±6.9 75.4±7.0 61.0±7.8 Retrain 51.0±1.4 33.5±4.5 44.0±5.5 32.0±5.2 1.62e11 (1.0×) 1.35e15 (1.0×) 4.49e07 (1.0×) FedEraser 51.2±0.5 35.7 (2.2±1.6) 49.1 (3.4±2.0) 47.4 (3.4±1.5) 1.62e11 (1.0×) 6.75e14 (2.0×) 8.98e10 (0.0005×) PGA 51.4±0.6 37.4 (4.3±4.1) 49.2 (6.4±4.9) 36.6 (5.3±5.3) 5.74e09 (28.2×) 5.10e13 (26.5×) 4.94e08 (0.09×) MoDe 50.9±1.1 41.9 (8.4±3.8) 55.0 (10.6±4.0) 42.3 (10.3±3.8) 2.19e10 (7.4×) 1.83e14 (7.4×) 8.98e07 (0.5×) FedAU 51.2±1.2 38.4 (4.5±2.2) 51.7 (6.7±2.5) 38.3 (5.3±1.9) 1.29e10 (12.5×) 1.08e14 (12.5×) 4.49e07 (1.0×) NoT 51.3±0.2 43.3 (9.9±4.4) 55.7 (11.7±5.7) 44.3 (12.3±5.5) 1.03e10 (15.7×) 8.64e13 (15.6×) 4.49e07 (1.0×) PUF-Special 52.3±0.4 36.8 (3.3±2.2) 47.3 (3.4±2.8) 35.9 (3.9±2.6) 4.01e09 (40.4×) 3.56e13 (38.0×) 4.49e07 (1.0×) PUF-Regular 52.2±0.6 35.4 (2.0±2.0) 48.8 (4.8±2.4) 36.4 (4.4±2.3) 6.44e09 (25.1×) 5.59e13 (24.1×) 4.49e07 (1.0×) CIFAR-100, Non-IID, ResNet-18, EE=10 Original 50.3±0.0 60.8±6.6 73.4±6.6 61.3±7.5 Retrain 48.0±0.8 31.6±4.5 42.1±5.2 31.2±4.9 1.62e11 (1.0×) 1.35e16 (1.0×) 4.49e07 (1.0×) PGA 46.3±1.9 34.7 (3.1±3.9) 45.3 (3.2±4.0) 33.5 (2.3±3.7) 3.01e14 (44.9×) 3.35e13 (40.4×) 4.94e08 (0.09×) MoDe 46.9±2.3 35.3 (3.9±3.6) 45.4 (4.7±4.1) 34.0 (3.6±3.3) 2.10e10 (7.7×) 1.76e15 (7.7×) 8.98e07 (0.5×) FedAU 46.8±2.4 36.8 (4.5±3.7) 47.1 (6.1±3.6) 34.7 (2.8±3.5) 1.13e10 (14.3×) 9.45e14 (14.3×) 4.49e07 (1.0×) NoT 46.7±2.2 39.7 (8.1±3.6) 48.9 (6.8±4.4) 37.2 (6.0±3.9) 6.22e09 (26.0×) 5.20e14 (26.0×) 4.49e07 (1.0×) PUF-Special 47.8±1.5 34.0 (2.4±2.9) 43.7 (1.6±3.2) 32.4 (1.2±3.0) 2.90e09 (55.8×) 2.67e14 (50.7×) 4.49e07 (1.0×) PUF-Regular 47.6±1.6 33.3 (1.7±2.8) 44.5 (2.4±2.9) 32.7 (1.5±2.7) 4.85e09 (33.4×) 4.44e14 (30.4×) 4.49e07 (1.0×) CIFAR-10, Non-IID, ResNet-18, EE=1 Original 83.7±0.0 88.6±3.9 84.4±5.4 81.4±6.3 Retrain 83.5±1.6 81.1±8.0 73.1±13.7 72.5±10.8 1.62e11 (1.0×) 1.35e15 (1.0×) 4.49e07 (1.0×) PGA 84.0±1.0 84.3 (3.2±4.9) 78.5 (5.4±7.7) 77.0 (4.5±6.0) 9.38e09 (17.3×) 8.14e13 (16.6×) 4.94e08 (0.09×) MoDe 83.5±1.6 82.3 (2.0±1.7) 72.6 (2.6±1.8) 71.8 (1.5±1.0) 2.40e10 (6.7×) 2.01e14 (6.7×) 8.98e07 (0.5×) FedAU 83.8±1.4 85.2 (4.0±2.6) 76.9 (4.3±3.7) 73.2 (2.9±2.9) 5.17e09 (31.3×) 4.32e13 (31.2×) 4.49e07 (1.0×) NoT 83.8±1.2 84.3 (3.1±3.0) 78.0 (4.9±4.5) 75.3 (2.9±3.4) 1.37e10 (11.9×) 1.14e14 (11.8×) 4.49e07 (1.0×) PUF-Special 83.9±1.2 82.0 (0.9±1.4) 72.0 (1.4±1.7) 71.1 (2.1±2.0) 3.30e09 (49.1×) 2.79e13 (48.5×) 4.49e07 (1.0×) PUF-Regular 83.8±1.2 82.4 (1.3±1.6) 72.8 (2.2±2.0) 71.5 (2.4±1.8) 5.31e09 (30.5×) 4.49e13 (30.1×) 4.49e07 (1.0×) CIFAR-100, Non-IID, MiT-B0, EE=1 Original 75.0±0.0 84.3±5.6 77.0±8.6 74.0±6.4 Retrain 73.3±0.8 57.8±5.3 48.6±5.0 44.6±3.9 1.19e10 (1.0×) 3.82e15 (1.0×) 1.33e07 (1.0×) PGA 73.6±0.4 62.2 (4.4±3.3) 53.6 (5.0±3.7) 50.5 (6.9±3.2) 1.68e09 (7.1×) 5.70e14 (6.7×) 1.46e08 (0.09×) MoDe 73.4±0.8 62.9 (5.2±2.2) 54.7 (6.1±3.1) 50.3 (5.9±2.9) 5.55e09 (2.1×) 1.78e15 (2.2×) 2.66e07 (0.5×) FedAU 73.6±0.7 63.7 (5.9±1.9) 55.2 (5.9±2.3) 50.2 (5.6±1.3) 3.54e09 (3.4×) 1.13e15 (3.4×) 1.33e07 (1.0×) NoT 73.4±0.7 67.9 (10.1±4.1) 59.4 (10.1±6.6) 54.0 (9.4±5.4) 3.70e09 (3.2×) 1.19e15 (3.2×) 1.33e07 (1.0×) PUF-Special 73.7±0.5 58.2 (0.4±1.5) 47.2 (1.4±1.7) 45.1 (0.5±1.4) 2.65e09 (4.5×) 8.42e14 (4.5×) 1.33e07 (1.0×) PUF-Regular 73.6±0.6 59.3 (1.5±1.7) 48.4 (2.7±1.8) 45.8 (1.2±1.3) 3.92e09 (3.0×) 1.26e15 (3.0×) 1.33e07 (1.0×)
TABLE III: Performance of PUF and other baselines across different settings for client unlearning. Baseline results are expressed as mean metric value (mean $\Delta$ $\pm$ standard deviation), with $\Delta$ in parentheses representing the average absolute difference from the Retrain baseline. Lower $\Delta$ corresponds to better unlearning performances. For efficiency metrics, higher $\times$ (reduction over Retrain baseline) are better. The best-performing method is shown in bold. When FedEraser [16] ranks first, it is underlined, as its efficiency limitations make it impractical in realistic scenarios.
V-D Main Results
Client Unlearning. Table III reports the performance of PUF’s best configurations (hyper-parameter tuning in Sec. V-E) and other baselines across multiple datasets (CIFAR-100 and CIFAR-10), data distributions (IID and Non-IID), and model architectures (ResNet-18, MiT-B0) in the single-client unlearning scenario. Results are averaged over ten independent runs, each with a different target client. Across all settings, PUF attains the lowest or near-lowest $\Delta$ values in forget metrics (e.g., Forget Accuracy), meaning it most closely replicates the retraining gold standard, thereby achieving highly effective forgetting.
[…]
In CIFAR-100 (Non-IID, ResNet-18, EE=1), PUF-Regular achieves the best $\Delta$ Forget Accuracy (2.0%), ahead of FedAU (4.5%), MoDe (8.4%), and NoT (9.9%), while preserving $25\times$ efficiency gains over retraining. PUF-Special slightly sacrifices forgetting (3.3% $\Delta$) for even higher efficiency ($\approx 40\times$). PGA achieves higher efficiency ($\approx 28\times$) but worse forgetting (4.3% $\Delta$ Forget Accuracy).
In CIFAR-100 (Non-IID, ResNet-18, EE=10), PUF-Regular exhibits the lowest $\Delta$ Forget Accuracy (1.7%) with $33\times$ efficiency gains, while PUF-Special offers the highest efficiency ($\approx 56\times$) and competitive forgetting (2.4%). PGA is also efficient ($\approx 44\times$) but less accurate in forgetting (3.1%).
In CIFAR-10 (Non-IID, ResNet-18), PUF-Special reaches the best forgetting performance ($\Delta$ Forget Accuracy <<1%) with the highest efficiency ($\approx 49\times$), outperforming all baselines in both dimensions.
Finally, in CIFAR-100 (Non-IID, MiT-B0), PUF-Special attains near-perfect forgetting ($\Delta$=0.4%) with solid efficiency ($\approx 4.5\times$), while PGA is the most efficient ($\approx 7\times$) but exhibits notably weaker forgetting ($\Delta$=4.4%). Overall, PUF variants consistently achieve an effective trade-off between efficacy and efficiency.
@@ V-D Main Results
Figure 4: Evolution of generalization ability (test accuracy) and forgetting effectiveness (forget accuracy) during the recovery phase across three different settings. Each pair of images presents Test Accuracy (Left) and Forget Accuracy (Right) for a representative client in a specific setting indicated in subcaption as a triple dataset, data distribution, model architecture. For our method, the charts only report the performance of PUF-Special for better visualization. For MoDe we do not show the multi-round unlearning phase for clarity.
Performance Consistency During Recovery. Figure 4 illustrates the evolution of Test Accuracy and Forget Accuracy throughout the recovery phase, beginning at recovery round 0, which corresponds to the application of the unlearning routine (either PGA, MoDe, FedAU, NoT, or our proposed solution, PUF). We do not report the evolution of the considered MIAs, as they closely follow the trends of Forget Accuracy. The figure presents results for three different settings, as indicated in the subcaptions. For reference, we also include values for both the Original Model and the Retrained Model. We emphasize the following key aspects for the full and easy understanding of the reported charts:
@@
Thus, Figure 4 provides valuable insights into the dynamics of these metrics during the recovery phase. The analysis highlights that PUF produces a superior unlearned model, significantly narrowing the gap with the retrained model earlier in the recovery phase compared to the other baselines. The figure clearly shows that the other baselines fail to achieve accurate forgetting by the end of the recovery phase, as evidenced by their larger gaps in Forget Accuracy. All the methods show an accuracy drop after unlearning. This is an inherent consequence of strategies that operate directly on the global model, as the removal of a client’s contributions inevitably results in a temporary degradation of overall performance [13].
Setting Method Efficacy Efficiency Test Acc. Forget Acc. ($\Delta$ $\downarrow$) MIA[Song] ($\Delta$ $\downarrow$) MIA[Yeom] ($\Delta$ $\downarrow$) Communication Computation Storage Bytes ($\times$ $\uparrow$) FLOPs ($\times$ $\uparrow$) Bytes ($\times$ $\uparrow$) CIFAR-100, IID, ResNet-18, EE=1 Retrain 59.4±0.4 58.5±0.5 56.0±0.6 49.2±0.7 1.80e11 (1.0×) 1.42e15 (1.0×) 4.49e07 (1.0×) FedAU 59.8±0.3 65.7 (7.2±1.1) 61.8 (5.8±0.9) 54.9 (5.7±0.9) 8.08e08 (222.7×) 7.12e12 (199.3×) 4.49e07 (1.0×) NoT 59.7±0.5 72.4 (13.9±0.7) 68.1 (12.1±1.0) 60.7 (11.5±1.0) 5.65e09 (31.8×) 4.99e13 (28.5×) 4.49e07 (1.0×) PUF-Special 60.0±0.8 64.5 (6.0±1.1) 60.6 (4.6±0.9) 53.8 (4.6±0.9) 1.43e09 (125.8×) 1.22e13 (116.4×) 4.49e07 (1.0×) PUF-Regular 60.1±0.8 64.6 (6.1±1.2) 60.7 (4.7±0.9) 53.9 (4.7±0.9) 2.51e09 (71.6×) 2.18e13 (65.3×) 4.49e07 (1.0×) CIFAR-100, Non-IID, ResNet-18, EE=1 Retrain 49.0±0.6 40.3±2.4 38.9±2.0 34.1±1.8 1.80e11 (1.0×) 1.42e15 (1.0×) 4.49e07 (1.0×) FedAU 50.0±0.2 47.2 (6.9±1.2) 45.1 (6.2±1.1) 40.2 (6.1±1.1) 8.08e08 (222.7×) 7.12e12 (199.3×) 4.49e07 (1.0×) NoT 49.6±0.8 49.0 (8.8±1.4) 47.0 (8.1±1.3) 41.8 (7.7±1.2) 6.46e09 (27.9×) 5.70e13 (24.9×) 4.49e07 (1.0×) PUF-Special 50.1±1.0 42.5 (2.2±1.2) 40.9 (2.0±1.1) 36.3 (2.2±1.1) 8.98e08 (200.5×) 7.50e12 (189.3×) 4.49e07 (1.0×) PUF-Regular 50.2±1.0 41.3 (4.3±1.3) 40.1 (4.0±1.2) 35.2 (4.1±1.2) 1.71e09 (105.5×) 1.46e13 (97.1×) 4.49e07 (1.0×)
TABLE IV: Performance of PUF and other baselines across different settings for sample unlearning when half of the dataset is the forget data. For efficacy metrics, baseline results are expressed as mean metric value (mean $\Delta$ $\pm$ standard deviation), with $\Delta$ in parentheses representing the average absolute difference from the Retrain baseline. Lower $\Delta$ corresponds to better unlearning performances. For efficiency metrics, higher $\times$ (reduction over Retrain baseline) are better. The best-performing method is shown in bold.